OpenStudy (anonymous):

What is the value of 134 + 2n^2 when n = 6? 68 98 170 206

OpenStudy (anonymous):

help @campbell_st

OpenStudy (campbell_st):

so you replace n with 6 which becomes \[134 + 2 \times 6^2 = \] just calculate it out

OpenStudy (anonymous):

but it say 4896

OpenStudy (campbell_st):

wow... you may like to recheck it...

OpenStudy (campbell_st):

6^2 = 36 so double it.... and then add 134

OpenStudy (anonymous):

ohh okay